OpenStudy (anonymous):

Compare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 − 4x y = −2x2 + 8x − 12 Look i need the answer ! this is my last question !

OpenStudy (anonymous):

Equation 1: y = x2 - 4x 1. concave up. [leading coefficient is positive] 2. vertex at (2,-4) [rewrite in canonical form as y=(x-h)^2+k =(x-2)^2-4, then (h,k) is the vertex.] 3. y-intercept = 0 [from the general form ax^2+bx+c, where c=0 => y-intercept=0] 4. zeroes at x=0 and x=4. [rewrite y=x^2-4x=x(x-4) => zeroes at x=0 and x=4].

OpenStudy (anonymous):

wt 2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

loll ok thanks ^,^